OpenStudy (anonymous):

Anyone know if there's a function in Mathematica that will give numerical answers to the inequality: \[2^n

OpenStudy (anonymous):

Could you explain a bit more with one example ?

OpenStudy (anonymous):

Actually, I only really need a function to solve: \[2^n = n!\] I'm looking for something similar to the reduce function. (Reduce doesn't seem to like factorials.) For example I can use Reduce[(x + 1)/(x + 2) < (x + 3)/(x + 4)] and Mathematica gives the output x<-4||x>-2. Does that make more sense?

OpenStudy (jamesj):

Well, just try some values here n = 1: 2^n = 2 > 1 = 1! n = 2: 2^n = 4 > 2 = 2! n = 3: 2^n = 8 > 6 = 3! n = 4: 2^n = 16 < 24 = 4! n = 5: 2^n = 32 < 120=5! and it's not hard to convince yourself that for all n > 5, 2^n < n!

OpenStudy (anonymous):

I understand that it is, and I know it's easy to prove by induction, but unfortunately I have to find a solution specifically through Mathematica and have been searching for quite a while.

OpenStudy (anonymous):

But the equation \[ 2^n =n! \] is only solvable for n = 0,and \[ 2^n \gt n! \] for \[ n \in [1,3]\], and after that \[ 2^n \lt n! \] for any \[ n \ge 4 \]isn't ?

OpenStudy (jamesj):

I see; I'm sorry, I don't know Mathematica commands.

OpenStudy (anonymous):

Thanks for the help anyway.

OpenStudy (anonymous):

Well then you have to write this specific function yourself,since I am not aware of any function that could do it by default.

OpenStudy (anonymous):

Ok, I was hoping one existed. Thanks for the help guys.

OpenStudy (zarkon):

are you supposed to use the gamma function? that is the only way to solve n!=2^n

OpenStudy (jamesj):

Yes, was thinking that too, but I assumed that because he was using n the values are integral.

OpenStudy (anonymous):

I doubt it. It's only undergraduate level and I think it would be overcomplicating it. And solutions are over natural n.

OpenStudy (jamesj):

For the record, the Gamma function is very much an undergraduate concept, but if n is integral, it's integral.

OpenStudy (anonymous):

Oh ok, I assumed it was a higher level probably being a tad naive.